OpenStudy (anonymous):

Solve the equation. x –2/7 = 3 @itsbribro i got 40%

OpenStudy (anonymous):

@itsbribro

OpenStudy (anonymous):

@itsbribro

OpenStudy (itsbribro):

ohh im sorry :(

OpenStudy (itsbribro):

http://web2.0calc.com/

OpenStudy (anonymous):

@itsbribro

OpenStudy (anonymous):

it is not working @itsbribro

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@nylearns

OpenStudy (itsbribro):

x=23/7

OpenStudy (anonymous):

yeah that is what i got but my test says something else

OpenStudy (itsbribro):

what does your test say?

OpenStudy (anonymous):

A. x = –3 2/7 B. x = –2 5/7 C. x = 2 5/7 D. x = 3 2/7

OpenStudy (anonymous):

@itsbribro Unit 5 | Lesson 12 Lesson Assessment: Equations with Fractions Quiz

OpenStudy (itsbribro):

hmm

OpenStudy (anonymous):

Solve. Express the answer in simplest form. 1/5+w=7 1/3

OpenStudy (anonymous):

@itsbribro

OpenStudy (anonymous):

Unit 5 | Lesson 12 Lesson Assessment: have u done this yet

OpenStudy (anonymous):

do u live in oakland CA

OpenStudy (itsbribro):

no i dont live there lol

OpenStudy (itsbribro):

i live in Roswell NM

OpenStudy (anonymous):

o wow

OpenStudy (anonymous):

Express the answer in simplest form. 1/5+w=7 1/3